OpenStudy (anonymous):

Find the sum of (-3)^(n-1) divided by 4^n. This is a geom. series. When I simplify it, im confused on what my "a_1" should be.

OpenStudy (anonymous):

\[\frac{(-3)^{n-1}}{4^{n}} = \]

OpenStudy (anonymous):

all I can see is that, by simplifying, you can get (-1/3) * (-3/4)^n. Does that help in any way?

OpenStudy (anonymous):

I have that too!!! Im going over an old test, and i got that answer wrong. The teacher said that a_1 should "1/4". But now that Im looking at it, i think it should be "1/3".

OpenStudy (anonymous):

1/4? Hmm... it could possibly be 4 as well I guess... 4 * (-3/4)^(n-1) but 1/4 doesn't look right

OpenStudy (anonymous):

I'm no expert tho... Maybe someone else could offer a second opinion?

OpenStudy (anonymous):

Yeaah, im wavering between the two. I agreed with 1/4 at first. But it looks weird now.

OpenStudy (anonymous):

The sum w/ (1/3) gives me 1.333 and the sum w/ (1/4) just gies me 1. Lol, they both kinda converge

OpenStudy (anonymous):

Series are the part of the semester which I've slept through... It's amazing I had even passed that course in the first place, so how do you determine when a serie converges? (if it's not too long to explain)

OpenStudy (anonymous):

My test on this stuff is today, so i HOPE i can even explain! If |r|<1, then it converges to Sum= a_1/1-r. If |r| >= 1 then the above geometric series diverges.

OpenStudy (anonymous):

ohhhhh, hold on, 1/4 is good too.... dunno why I screwed up that one, but you can get (-1/3)(-3/4)^n or (1/4)(-3/4)^(n-1)... So 4 * (-3/4)^(n-1) is WRONG and I do apologize. On your bottom term, 4^n = 4(4^(n-1))... I dunno why you would pick to have your exponent as n-1 instead of n, but I hope you do ;-)

OpenStudy (anonymous):

Yes, you are pulling out the "4", which means you are subtracting. thats why you have the "n-1". Doing so makes the exponents equal. so (-3^n-1)/(4^n-1). You see? lol

OpenStudy (anonymous):

yeah, I guess, but I would naturally prefer pulling out the -3 from the top bit, to make the exponent n. It just looks cleaner with n, but it's probably not how to do solve a geometric series... anyways, I think I've caused more confusion than anything here (at least, in my brain I did)

OpenStudy (anonymous):

THIS is why I dont like math. Thanks anyways. Since you had to take this course, what are you majoring in??

OpenStudy (anonymous):

oh, it's been at least 10 years. I live in Quebec and this course was part of what we call a "pre-university" degree (it's a weird education system here), taken in college. I then left math behind to try my hand at biology, which was even more boring ;-)

OpenStudy (anonymous):

Some schools here in Georgia & US have that. Im doing electrical engineering and im bored as well. Thanks for your help!!!!!!!!

OpenStudy (anonymous):

I'm not sure I'm doing this correctly, but I seem to get that this serie converges to 1/7... I'm doing this : s = 1/4 - 3/16 +9/64 - 27/256 + ... + 1/4(-3/4)^(n-1) -3s/4 = -3/16 + 9/64 - 27/256 + ... + 1/4(-3/4)^n s-(-3/s4) = 1/4 + 1/4(-3/4)^n 7s/4 = 1/4(1 + (-3/4)^n) s = 1/7 ( 1 +(-3/4)^n) if n tends to infinity, s = 1/7, right?

OpenStudy (anonymous):

oh, when you said you had gotten 1 with 1/4, I think you might've done s - 3/4s, correct? s/4 = 1/4(1+(-3/4)^n) s = 1+(-3/4)^n s = 1

OpenStudy (anonymous):

Its righte, but You definintely did it the hard way. S= (1/4) / (1-(3/4)) Use formula: 1st term / (1- |r|)

OpenStudy (anonymous):

are you sure that you need the absolute value of r in your formula? because Wolfram seems to tell that this series converges to 1/7 as well, and using absolute value doesn't make sense mathematically :-S

OpenStudy (anonymous):

Hey, thats what the book and the teacher says. I have learned no other way. Because we have r= (-3/4), So |r|=|-3/4|= 3/4 <1, it converges. Then use the sum thingy

OpenStudy (anonymous):

oh, yes, to determine that it converges, you definitely need to use the absolute value, but to determine the actual value towards which it converges, I think that you don't need it (at least, I couldn't find it in google, but I did find instance of s = a_1/(1-r) tho)

OpenStudy (anonymous):

and that's why I hate formulas, because if you remember them wrong or if you learnt them wrong, you can be wrong... @experimentX @foolformath guys (or girls, I dunno), can you help us out here please, or summon people which you know could help?

OpenStudy (anonymous):

Yes, I said that. I never said that you need |r| in the sum thingy. You only need |r| to see if it converges.

OpenStudy (experimentx):

(-3)^(n-1) / 4^n = (-1)^(n-1)*(3/4)^n/3 <--- looks easier now.

OpenStudy (experimentx):

both are definitely guys @m_charron2

OpenStudy (anonymous):

\[ a_1=\frac 1 4 \]

OpenStudy (experimentx):

\( \frac{(-1)^{n-1}}{3} (\frac{3}{4})^n\) put the values of n, you should be easily to find.

OpenStudy (experimentx):

a1 = 1/4 a2 = -3/16 a3 = +9/64 .. ..

OpenStudy (anonymous):

I will like to represent the n-th term as:\[a_n = \frac 14 \times \left(-\frac 34\right)^{n-1} \]

OpenStudy (anonymous):

now \(|r| = \frac 3 4 \lt 1 \), hence the series converges.

OpenStudy (anonymous):

Thanks everyone!